One object is moving and one object is at rest. Th
Answer #1

Given m1= 5 kg and m2 = 8.5 kg

Before collision v_1i = 22.5 m/s and v_2i = 0

From conservation of momentum m1 v_1i + m2 v_2i = (m1 + m2) v_f

                            v_f = 5 * 22.5 /( 5+8.5) = 8.33 m/s

K.E before collision = 1/2 * 5 * (22.5)2 = 1265.6 J

K.E after collision = 1/2 (5+8.5) * (8.33)2 = 468.37 J